# Improving and Evaluating Open Deep Research Agents

### v1

@@ Abstract
We focus here on Deep Research Agents (DRAs), which are systems that can take a natural language prompt from a user, and then autonomously search for, and utilize, internet-based content to address the prompt. Recent DRAs have demonstrated impressive capabilities on public benchmarks however, recent research largely involves proprietary closed-source systems. At the time of this work, we only found one open-source DRA, termed Open Deep Research (ODR). In this work we adapt the challenging recent BrowseComp benchmark to compare ODR to existing proprietary systems. We propose BrowseComp-Small (BC-Small), comprising a subset of BrowseComp, as a more computationally-tractable DRA benchmark for academic labs. We benchmark ODR and two other proprietary systems on BC-Small: one system from Anthropic and one system from Google. We find that all three systems achieve 0% accuracy on the test set of 60 questions. We introduce three strategic improvements to ODR, resulting in the ODR+ model, which achieves a state-of-the-art 10% success rate on BC-Small among both closed-source and open-source systems. We report ablation studies indicating that all three of our improvements contributed to the success of ODR+.
@@ 1  Introduction
In this work we focus on the problem of developing Deep Research Agents (DRAs), wherein our goal is to develop a system that can take as input a natural language prompt from a user, then autonomously search for and utilize internet-based content to address the prompt. This is a challenging problem because, in principle, it typically comprises several sub-problems that are each challenging for contemporary artificial intelligence methods: for example, breaking a natural language prompt into easier sub-questions, reasoning about the use of an internet search engine to find relevant information on the internet, and then reasoning about that retrieved content to address the original prompt. Until recently there has been relatively limited progress in this area due to the difficulty of the problem, however large language models (LLMs) have demonstrated the potential to address many of these challenges and several organizations have developed proprietary systems that seek to perform Deep Research. Examples include OpenAI’s recent Deep Research [3] and Google’s Deep Research [6], Perplexity’s research capabilities [2].
@@
One challenge with LLM-based DRAs is evaluating their performance. Benchmark problems for DRAs should simultaneously satisfy two competing properties. First, the problems must ideally be sufficiently challenging so that they cannot be easily addressed by existing methods, such as a single prompt to a LLM, or a simple single query to a browser (made either by a human or an LLM). Given the impressive capabilities of LLMs in recent years, the difficulty of problems that are suitable for DRAs has increased rapidly. Some existing benchmarks are potentially suitable for DRAs, such as HotpotQA [5] and Natural Questions [17], however recent LLM-based methods have achieved near-perfect accuracy, motivating the need for more challenging benchmark problems. Building such a benchmark is challenging however, because any benchmark questions should also include ground truth solutions, and also importantly, it should be possible to find these solutions on the internet. Without this latter property, the benchmark problems cannot be solved by a DRA. Therefore DRA benchmark problems must simultaneously be so difficult that their solutions are difficult to find, but we must also be certain that there is a solution that can be found on the internet.
@@
Very recently, the BrowseComp [4] benchmark was introduced to address the limitations of existing benchmarks. BrowseComp includes over 1200 problems that are, by design, challenging to solve both for humans and existing LLM-based systems, while also being highly likely to have solutions that can be found via internet search. The authors of BrowseComp [4] benchmarked several proprietary systems, finding that most existing DRAs performed poorly, achieving less than 10% accuracy. The very best system, which utilized specialized methods, and substantial test-time compute, was able to achieve 50% accuracy. One major limitation of BrowseComp however is that it focused entirely on proprietary closed-source DRAs from OpenAI, creating a limited understanding of existing DRA capabilities, and a significant barrier for researchers seeking to understand, evaluate, and advance DRAs. At the time of this work, there was only one open-source DRA, termed Open Deep Research (ODR) [8]. However, the performance of ODR has yet to be quantitatively evaluated, making it unclear how open source DRAs compare with proprietary counterparts, and there is currently no baseline methods upon which to improve DRAs within the open research community.
@@
Contributions of this work. To address these limitations, we develop BrowseComp-Small, a more computationally tractable deep research benchmark, comprising two disjoint sets of sixty questions sampled from BrowseComp: a training set, intended for DRA development; and a testing set, intended for DRA performance evaluation. We found that ODR is unable to answer any of the challenging questions in the BrowseComp-Small testing set. We propose several methodological improvements to the ODR system to support more effective deep research, resulting in our proposed ODR+ system. ODR+ successfully answers 20% of the training and 10% of the test BrowseComp questions, and therefore greatly outperforms the original ODR system. We also find that ODR+ outperforms several proprietary closed-source systems as well. Using ablation studies on the BrowseComp benchmark, we demonstrate the benefit of each of our proposed methodological improvements to ODR. We provide an open-source implementation of ODR+ [^1] https://github.com/EngDoLabadi/odrplus. We summarize our contributions as follows:
@@ 2  Related Work
Deep Research Agents. Early work in autonomous deep research began with WebGPT (2021) [1], which allowed LLMs to issue real Bing search queries and cite results. However, it supported only single-turn QA and was not released as a reusable browsing agent, making it unsuitable for benchmarking here. Recent proprietary systems represent a major advance in apparent capability, which has recently been substantiated through some empirical benchmarking (discussed below). Notable examples include OpenAI’s Deep Research, Google Gemini Deep Research, and Perplexity AI, which demonstrate varying degrees of real-time search and multi-step information retrieval. Despite their impressive performance however, these systems are closed-source, motivating the need for open, analyzable alternatives.
@@
DeepResearcher [14] introduced a reinforcement learning framework for training browsing agents that autonomously decide what to search, read, and extract from the web. Another recent system is WebThinker [13], proposed a modular architecture to interleave deep Web exploration with reasoning, focusing on scientific and factual question answering. Although both provide code bases, their implementations are complex and are not easy to reuse or modify. Due to our limited computational budget and their very recent introduction, we did not include these methods in our benchmarking.
Existing Deep Research Benchmarks. Multi-hop reasoning across sources is a key challenge for Deep Research agents. Benchmarks such as HotpotQA [5] and 2WikiMultiHopQA [9] focus on 2-hop questions but operate within a closed setting, where the agent must find answers using a fixed set of Wikipedia passages. These benchmarks do not require open-ended web search or the ability to generate and revise search queries. As a result, they do not test whether systems can autonomously generate effective queries, plan multi-step retrieval strategies, or identify when additional information is needed, capabilities that are central to Deep Research agents. By contrast, the recent BrowseComp benchmark [4] is explicitly designed to evaluate complex, multi-hop question answering that requires synthesizing information from multiple web sources. BrowseComp questions require issuing multiple web searches, retrieving diverse pieces of evidence, and synthesizing information across documents, making it a more appropriate benchmark for studying Deep Research agents.
In addition to BrowseComp, other recent benchmarks have further advanced the evaluation and design of Deep Research agents. Mind2Web 2 [11] introduces 130 long-horizon tasks that require agents to browse unfamiliar websites and generate structured, cited answers. It also proposes an ”agent-as-a-judge” framework where agents evaluate the factuality of their own outputs. However, the benchmark emphasizes self-assessment rather than complex information synthesis and was released too recently for inclusion in our study.
Deep Research Bench [12] focuses on 89 live web tasks that test retrieval and synthesis across open-ended queries. It provides reference answers and clear evaluation criteria to assess information accuracy and completeness, but it covers fewer domains and question types than BrowseComp, with an emphasis on factual verification and summarization rather than multi-hop or exploratory reasoning. Huang et al. [10] provide a broad taxonomy and analysis of Deep Research systems, offering a high-level overview of how recent benchmarks and agent designs relate to one another.
@@ 3  The BrowseComp-Small Benchmark
We first introduce the BrowseComp (BC) benchmark, from which we create the BrowseComp-Small (BC-Small) benchmark that we use in our experiments to evaluate competing Deep Research systems. BrowseComp. [4] is a benchmark for Deep Research agents comprising 1,266 questions covering a wide range of topics, including entertainment, science, history, politics, and geography. BrowseComp questions were constructed so that they are challenging to solve, but to have short answers that are easy to verify. Questions were constructed by human ”trainers” using the following procedure. Each question was constructed by first identifying some object (e.g., a person, place, or thing), and then selecting a set of properties about that object that would, collectively, uniquely identify it. Using the selected set of properties, the designer would construct a question that asks the DRA to identify the object that satisfies all of the selected properties. For example, one BrowseComp question asks: ”Which 90s TV series starred an actor born in Tennessee, an actor who was a Caribbean immigrant, and an actor whose father was a law enforcement officer for more than 3 decades? The series was short-lived.”. The authors of BrowseComp used various criteria to ensure the difficulty of each question. For example, human trainers ensured that each question could not be correctly answered by another person within ten minutes. They also confirmed that existing models such as ChatGPT (with and without browsing) and an early version of OpenAI’s Deep Research were unable to solve them. See [4] for full design details.
BrowseComp-Small (BC-Small) The computational cost of evaluating DRAs on the full BrowseComp benchmark is large, especially for academic labs. To make BrowseComp more accessible while maintaining its utility as a benchmark, we created a smaller benchmark - termed BrowseComp-Small (BC-Small) - that comprises a subset of 120 questions from BrowseComp. We selected the questions to maintain a similar distribution of topics. Crucially, and in contrast to BrowseComp, we split BC-Small into 60 questions that are used for DRA development - essentially a training set - and another disjoint 60 questions as a validation set. This should help ensure generalization. Our choice of 120 questions was chosen to be consistent with the size of other recent public benchmarks, such as Deep Research Bench [12] (89 questions) and Mind2Web2 [11] (130 questions), where the authors also cited the high cost of issuing multiple search queries, parsing content, and performing iterative reasoning for each example.
@@ 4  Open Deep Research
Here we describe the ODR system from [8], upon which our proposed ODR+ is based. Here we provide the essential system-level details of its operation. An open-source implementation is also available [^2] ODR Implementation: https://github.com/nickscamara/open-deep-research..
The operation of ODR is illustrated in Fig. 1 and consists of three main steps.
(1) Web URL Search. The user submits a natural language question, which is passed to a large language model (LLM) along with a system prompt instructing it to generate a concise search query suitable for an internet search engine. The LLM produces a single general-purpose query without decomposing the question or performing intermediate reasoning.
(2) Content Extraction. ODR submits the generated query to a search engine and retrieves a list of candidate web pages. It opens the top-ranked link and extracts the rendered (i.e., human-visible) content of the page, which is converted to plain text and passed forward without additional filtering or parsing.
(3) Response Generation. The extracted text and original user question are then passed to the LLM, along with a prompt instructing it to generate an answer using only the retrieved content. The model produces a free-form response—a natural language answer that is not constrained to a fixed structure—which is returned to the user.
@@ 5  Open Deep Research Plus (ODR+)
Here we describe our system, ODR+, which is constructed by making several improvements to the ODR system. ODR provides an important initial working system, but it suffers from several limitations that cause it to fail on complex, multi-hop research questions. We hypothesize that ODR fails on these problems for at least three reasons: it does not decompose the user query into simpler sub-questions; it lacks any form of iterative reasoning or adaptive planning; and it is not prompted to produce structured output. ODR+ addresses these limitations through the introduction of three modules illustrated in Fig 2: Question Decomposition, Sub-solution Search, and Response Synthesis. We next describe each of these three major modules, as well as sub-modules that contribute to them, which are detailed in pseudocode for ODR+ in Algorithm 1. The engineered prompts used throughout the ODR+ pseudocode are summarized in Table 1.
@@
$P_{1}$ Constraint Extraction Ask the LLM to extract the most specific constraints (e.g., names, dates, descriptors) from the user query to guide downstream reasoning.
@@
Table 1: Summary of engineered prompts used in ODR+ system, ordered by execution sequence.
Question Decomposition
@@
The first module of ODR+ converts the original user query, $userQuery$, into a set of focused sub-questions, as detailed in lines 7-8 of Algorithm 1. This begins with a call to extractConstraints (Line 7), which takes as input the prompt $P_{1}$ in Table 1 and the original query. The prompt instructs the language model to extract highly specific identifying constraints—such as names, dates, locations, and numerical values—that help narrow the answer space. The output is a structured JSON array of constraints.
@@
Next, the system calls generateSubQuestions (Line 8), which receives $P_{2}$, the original query, and the extracted constraints. This prompt guides the model to generate a small number of clear, fact-based sub-questions that target the extracted constraints. The resulting sub-questions are stored in the queue S.subquestions, which forms part of the system’s internal research state S. This state also tracks retrieved evidence, depth of search, processed URLs, and intermediate results, as initialized in Lines 4–5.
Iterative Sub-Solution Search

### v2

@@ Abstract
Deep Research Agents (DRAs) are systems that can take a natural language prompt from a user, and then autonomously search for, and utilize, internet-based content to address the prompt. Recent DRAs have demonstrated impressive capabilities on public benchmarks. However, recent research largely focuses on proprietary closed-source systems. At the time of this work, we identified only one open-source DRA, termed Open Deep Research (ODR). In this work, we adapt BrowseComp, the challenging recent benchmark dataset, to compare ODR to existing proprietary systems. We propose BrowseComp-Small (BC-Small), comprising a subset of BrowseComp, as a more computationally-tractable DRA benchmark for academic labs. We benchmark ODR and two other proprietary systems on BC-Small: one system from Anthropic and one system from Google. We find that all three systems achieve 0% accuracy on the test set of 60 questions. We introduce three strategic improvements to ODR, resulting in the ODR+ model, which achieves a state-of-the-art 10% success rate on BC-Small among both closed-source and open-source systems. We report ablation studies indicating that all three of our improvements contributed to the success of ODR+.
@@ 1 Introduction
In this work, we focus on the problem of developing Deep Research Agents (DRAs), wherein our goal is to develop a system that can take as input a natural language prompt from a user, then autonomously search for and utilize internet-based content to address the prompt. This is a challenging problem because, in principle, it typically comprises several sub-problems that are each challenging for contemporary artificial intelligence methods: for example, breaking a natural language prompt into easier sub-questions, reasoning about the use of an internet search engine to find relevant information on the internet, and then reasoning about that retrieved content to address the original prompt. Recently however, large language models (LLMs) have demonstrated the potential to address many of these challenges and several organizations have developed proprietary systems that seek to perform Deep Research. Examples include OpenAI’s recent Deep Research [11], Google’s Deep Research [2], and Perplexity’s research capabilities [12].
@@
One challenge with LLM-based DRAs is evaluating their performance, because the problems should, ideally, simultaneously satisfy two major competing properties. First, the problems must be sufficiently challenging so that they cannot already be easily solved by existing methods, such as a single prompt to an LLM, or a simple single query to a browser. Some existing benchmarks are theoretically suitable for DRAs, such as HotpotQA [13] and Natural Questions [7], however recent LLM-based methods have achieved near-perfect accuracy, motivating the need for more challenging benchmarks. The increasing difficulty of the benchmark problems however makes it difficult to satisfy the 2nd needed property: any benchmark question should also include ground truth solutions to enable performance evaluation. Furthermore, it should be possible to find these solutions on the internet, or else they cannot be solved by a DRA. Therefore DRA benchmark problems must simultaneously be so difficult that their solutions are difficult to find, even by a human, but we must also be certain that there is a solution, and that it can be found on the internet.
@@
Very recently, the BrowseComp [10] benchmark was introduced to address the limitations of existing benchmarks. BrowseComp includes over 1200 problems that are, by design, challenging to solve both for humans and existing LLM-based systems, while also being highly likely to have solutions that can be found via internet search. The authors of BrowseComp [10] benchmarked several proprietary systems from OpenAI, and found that all of the systems (except one) performed poorly, achieving less than 10% accuracy. The very best system, which utilized specialized methods, and substantial test-time compute, was able to achieve 50% accuracy. One major limitation of the existing evaluation of BrowseComp is that it has so far focused entirely on proprietary, closed-source DRAs from OpenAI. This creates a limited picture of DRA capabilities, as open-source systems have not yet been systematically benchmarked. In practice, this is due in part to the high computational cost of running BrowseComp at scale, which has so far restricted thorough evaluation to organizations with access to substantial computational resources. At the time of this work, there was only one open-source DRA, termed Open Deep Research (ODR) [1]. However, the performance of ODR has yet to be quantitatively evaluated, making it unclear how open source DRAs compare with proprietary counterparts, and there are currently no baseline methods upon which to improve DRAs within the open research community.
@@
Contributions of this work. To address these limitations, we propose BrowseComp-Small, a more computationally tractable deep research benchmark, comprising two disjoint sets of sixty questions sampled from BrowseComp: a training set, intended for DRA development; and a testing set, intended for DRA performance evaluation. We found that ODR is unable to answer any of the challenging questions in the BrowseComp-Small testing set. We propose several methodological improvements to the ODR system to support more effective deep research, resulting in our proposed ODR+ system. ODR+ successfully answers 20% of the training and 10% of the test BrowseComp questions, and, therefore, greatly outperforms the original ODR system. We also find that ODR+ outperforms several proprietary closed-source systems as well. Using ablation studies on the BrowseComp benchmark, we demonstrate the benefit of each of our proposed methodological improvements to ODR. We provide an open-source implementation of ODR+  [^1] https://github.com/EngDoLabadi/odrplus. We summarize our contributions as follows:
@@ 2 Related Work
Deep Research Agents. Early work in autonomous deep research began with WebGPT (2021) [9] was the first piece of autonomous deep research work that enabled LLMs to ask actual Bing search questions and cite the results. It was unsuitable for benchmarking here, though, because it only supported single-turn QA and was not made available as a reusable browsing agent. Recent empirical benchmarking has confirmed that recent proprietary systems represent a significant improvement in apparent capability (discussed below). Prominent instances of real-time search and multi-step information retrieval include Google Gemini Deep Research, Perplexity AI, and OpenAI’s Deep Research. The need for open, analyzable alternatives is prompted by the fact that these systems are closed-source despite their remarkable performance.
@@
DeepResearcher [14] introduced a reinforcement learning framework for training browsing agents that autonomously decide what to search, read, and extract from the web. Another recent system is WebThinker [8], proposed a modular architecture to interleave deep Web exploration with reasoning, focusing on scientific and factual question answering. Although both provide code bases, at the time of our experiments they were not straightforward to reproduce; DeepResearcher required reinforcement-learning training runs with substantial GPU resources, and WebThinker’s modular pipeline involved custom integration steps that were not yet fully documented or packaged for reuse. Given our limited compute budget and the fact that these systems had only just been released (e.g., and lacked full documentation) we were unable to obtain reliable runs suitable for benchmarking. We therefore restricted our comparisons to ODR, ODR+, and proprietary systems for which evaluation was feasible.
Existing Deep Research Standards. One of the main challenges for Deep Research agents is multi-hop reasoning across sources, where a hop is a single step of reasoning that links one piece of evidence to another or a question to a piece of evidence. HotpotQA [13] and 2WikiMultiHopQA [4] are benchmarks that focus on 2-hop questions, but they do so in a closed Wikipedia environment without open-ended search or query revision. Therefore, they do not adequately assess whether agents are capable of planning multi-step retrieval, generating queries on their own, or determining when additional evidence is required. The BrowseComp benchmark [10], on the other hand, is a better test for Deep Research agents because it is made for complex multi-hop QA, which calls for conducting numerous web searches, obtaining a variety of evidence, and combining information from different documents.
In addition to BrowseComp, other benchmarks have very recently been developed and advanced the evaluation of DRAs. These benchmarks were released too recently for us to consider or include in our work, however, we describe them here to account for important related progress in this fast-paced research area. Mind2Web2 [3] introduces a ”agent-as-a-judge” framework for self-assessment and consists of 130 long-horizon tasks that require agents to browse unknown websites and generate structured, cited answers. However, the benchmark was published too recently to be included in our study and places more emphasis on self-assessment than on complex information synthesis. Though it covers fewer domains than BrowseComp and prioritizes factual verification over multi-hop reasoning, Deep Research Bench [5] consists of 89 live web tasks with reference answers and explicit evaluation criteria. A taxonomy and analysis of recent DRA benchmarks and system designs are also provided by Huang et al. [6].
@@ 3 The BrowseComp-Small Benchmark
We first introduce the BrowseComp (BC) benchmark, from which we create the BrowseComp-Small (BC-Small) benchmark used in our experiments. BrowseComp. [10] is a benchmark for Deep Research agents comprising 1,266 questions on a wide range of topics, including entertainment, science, history, politics, and geography. BrowseComp questions were constructed so that they are challenging to solve, even for humans, but have short answers that are easy to verify. Questions were constructed by human ”trainers” using the following procedure. Each question was constructed by first identifying some object (e.g., a person, place, or thing), and then selecting a set of properties about that object that would, collectively, uniquely identify it. Using the selected set of properties, the designer would construct a question that asks the DRA to identify the object that satisfies all of the selected properties. For example, one BrowseComp question asks: ”Which 90s TV series starred an actor born in Tennessee, an actor who was a Caribbean immigrant, and an actor whose father was a law enforcement officer for more than 3 decades? The series was short-lived.”. The authors of BrowseComp used various criteria to ensure the difficulty of each question. For example, human trainers ensured that each question could not be correctly answered by another person within ten minutes. They also confirmed that existing models such as ChatGPT (with and without browsing) and an early version of OpenAI’s Deep Research were unable to solve them. See [10] for full design details.
BrowseComp-Small (BC-Small) The computational cost of evaluating DRAs on the full BrowseComp benchmark is large, especially for academic labs. To make BrowseComp more accessible while maintaining its utility as a benchmark, we created a smaller benchmark - termed BrowseComp-Small (BC-Small) - that comprises a subset of 120 questions from BrowseComp. We sampled questions to maintain a similar distribution of topics as the full BrowseComp benchmark. Crucially, and in contrast to BrowseComp, we split BC-Small into 60 questions that are used for DRA development - essentially a training set - and another disjoint 60 questions as a testing set, with the goal of better evaluating DRA generalization. Our choice of 120 questions was chosen to be consistent with the size of other recent public benchmarks, such as Deep Research Bench [5] (89 questions) and Mind2Web2 [3] (130 questions), where the authors also cited the high cost of issuing multiple search queries, parsing content, and performing iterative reasoning for each example.
@@ 4 Open Deep Research
Here we describe the ODR system from [1], upon which our proposed ODR+ is based. We provide the essential system-level details of its operation, but further implementation details can be found in the supplemental information. An open-source implementation of ODR is also available [^2] ODR Implementation: https://github.com/nickscamara/open-deep-research..
The operation of ODR is illustrated in Fig. 1 and consists of three main steps. (1) Web URL Search. The user submits a natural language question, which is passed to a large language model (LLM) along with a system prompt instructing it to generate a concise search query suitable for an internet search engine. The LLM produces a single general-purpose query without decomposing the question or performing more reasoning. (2) Content Extraction. ODR submits the generated query to a search engine and retrieves a list of candidate web pages. It opens the top-ranked link and gets the rendered (i.e., visible to humans) content of the page, which is then converted to plain text without any extra filtering or parsing. (3) Generating a response. The LLM gets the extracted text and the original user question, along with a prompt instructing it to use only the retrieved content to generate an answer. The model gives the user a free-form answer, which is a natural language answer that doesn’t have to follow a specific format.
@@ 5 Open Deep Research Plus (ODR+)
Here we describe our system, ODR+, which is constructed by making several improvements to the ODR system. ODR provides an important initial working system, but it suffers from several limitations that cause it to fail on complex, multi-hop research questions. We hypothesize that ODR fails on these problems for at least three reasons: it does not decompose the user query into simpler sub-questions; it lacks any form of iterative reasoning or adaptive planning; and it is not prompted to produce structured output. ODR+ addresses these limitations through the introduction of three modules illustrated in Fig 2: Question Decomposition, Sub-solution Search, and Response Synthesis. We next describe each of these three major modules, as well as sub-modules that contribute to them, which are detailed in pseudocode for ODR+ in Algorithm 1. The engineered prompts used in the ODR+ pseudocode are summarized in Table 1, and the full prompts are provided in the supplement.
@@
$P_{1}$ Constraint Extraction Asks the LLM to extract the specific constraints (e.g., names, dates, descriptors) from the user query to guide downstream reasoning.
@@
Table 1: Summary of engineered prompts used in ODR+ system, ordered by execution sequence. Full prompt text is included in the supplementary material.
5.1 Question Decomposition
@@
The first module of ODR+ converts the original user query, $userQuery$, into a set of focused sub-questions, as detailed in lines 7-8 of Algorithm 1. This begins with a call to extractConstraints (Line 7), which takes as input the prompt $P_{1}$ in Table 1 and $userQuery$. The prompt instructs the language model to extract explicit identifying details—such as names, dates, locations, or numerical values—that help narrow the search space. The output is returned in a simple structured format (e.g., a JSON list of constraints). For example, given the query ”Which 90s TV series starred an actor born in Tennessee and an actor who was a Caribbean immigrant?”, the model would extract constraints like ["1990s","actor born in Tennessee","Caribbean immigrant"].
@@
Next, the system calls generateSubQuestions (Line 8), which receives $P_{2}$, and $userQuery$, and the extracted constraints from $P_{1}$ above. Prompt $P_{2}$ guides the model to generate a small number of clear, fact-based sub-questions that target the extracted constraints. The resulting sub-questions are stored in the queue S.subquestions, which forms part of the system’s internal research state S. This state also tracks retrieved evidence, depth of search, processed URLs, and intermediate results, as initialized in Lines 4–5.
5.2 Iterative Sub-Solution Search